OpenStudy (anonymous):

2x+3y=4 4x+6y=9 using substitution

terenzreignz (terenzreignz):

You'll be in for a slight shock... :)

terenzreignz (terenzreignz):

Well, you know what they say, Herr Euler... great mind think...uhh. of ice cream... ANYWAY.... To use substitution, pick one of the variables in the first equation (x or y, in this case, it doesn't really matter) and isolate it.

OpenStudy (anonymous):

soo

OpenStudy (dan815):

(2x+3y=4)*2 -(4x+6y=9) --> -2x=-1 x=1/2 y=1

terenzreignz (terenzreignz):

What kind of sorcery is this? o.O

terenzreignz (terenzreignz):

x cancels out, @dan815 and so does y

OpenStudy (campbell_st):

wow and I thought it was 0 = -1